OpenStudy (anonymous):

mr jennings is choosing a comittee of 3 out of his class of 26. how many different ways can he choose his committe of 3

OpenStudy (dumbcow):

this uses combination's the answer is 26 choose 3 and the formula uses factorals 26 C 3 = 26!/(26-3)!*3! = 26*25*24*...1/(23*22*...1)*(3*2*1) things will then cancel, multiply whats left over and thats your answer

OpenStudy (anonymous):

thanks for youe name to be dumb cow your very smart. lol thanks helped alot

OpenStudy (anonymous):

wait i got 26 is that wrong?

OpenStudy (dumbcow):

you should end up with (26*25*24)/(3*2*1) = 15600/6 = 2600

OpenStudy (dumbcow):

most scientific and graphing calculators have a function for combinations so you can check your answers,